Determine the amount of energy needed to melt a 5.0 kg block of lead if it is heated from a temperature of 23.0 degrees Celsius. The melting point of lead is 327.5 degrees Celsius.
Answer #1

Given,

Mass of lead,  m 5kg

Initial temperature, T_i=23\degree C

Final Temperature, T_f=327.5\degree C

We know that specific heat of lead is 130Jkg^{-1}K^{-1} and latent heat of fusion of lead is 23000J/kg .

When we start heating the block, two thing will happen:-

1) The temperature will increase from initial to final temperature.

2) State of lead will change from solid to liquid.

Heat reqquired to change the temperature from T_i to T_f is,

Q_1 = ms\Delta T

Q_1 = 5*130*(327.5-23)

Q_1 = 197925J

Heat required to change the state of lead from solid to liquid is,

Q_2=m*l

Q_2=5*23000

Q_2=115000J

Total heat required, Q=Q_1+Q_2

Q=(197925+115000)J

Q=312925J

Q=3.13*10^{5}J